# Exact Quench Dynamics from Thermal Pure Quantum States

### v2

@@ Abstract
We present an exact solution for the real-time dynamics following a quench from a thermal pure quantum (TPQ) state in an integrable system. Although equilibration is expected in this setting because the TPQ state already encodes the thermal expectation values of all conserved quantities, the approach to equilibrium shows highly nontrivial coherent dynamics. In the spin-1/2 XX chain, local observables become stationary at their thermal values after dephasing, while the entanglement entropy exhibits a characteristic double-plateau structure. We obtain this behavior exactly using three complementary approaches: two-dimensional (2D) conformal field theory (CFT) on the Klein bottle, an exact numerical evolution based on the matrix Riccati equation, and an asymptotically exact quasiparticle picture. These results demonstrate that the non-monotonic entanglement evolution is a macroscopic manifestation of the coherent dephasing of anomalous pairing correlations in the initial TPQ state.
A central pillar of modern statistical mechanics is the eigenstate thermalization hypothesis (ETH), which provides a framework for understanding how isolated, chaotic quantum systems approach equilibrium [1, 2, 3, 4]. The ETH asserts that individual energy eigenstates of chaotic systems are already locally thermal and reflect canonical typicality, meaning that a generic pure state from a microcanonical energy shell is locally indistinguishable from a thermal ensemble [5]. Consequently, the real-time evolution from a state that is already thermal-like, such as a thermal pure quantum (TPQ) state [6], is expected to display little or no dynamical change in local observables or entanglement entropy. In contrast, integrable systems, endowed with an extensive number of conservation laws, do not follow ETH and retain detailed memory of their initial conditions, relaxing instead to nontrivial stationary states described by a generalized Gibbs ensemble (GGE) [7, 8].
This raises a fundamental question: how does an integrable system evolve if it is prepared in a TPQ state? Does the dynamics retain the imprint of integrability, or reflect the thermal nature of the initial state? In this Letter, we provide an exact solution to this question and uncover a third dynamical paradigm, which is neither the trivial evolution of a chaotic system nor the standard non-thermal relaxation to a GGE. We consider a special TPQ state $|\Psi_{\beta}\rangle$ in the integrable spin-1/2 XX chain, with both the preparation and time evolution governed by the same Hamiltonian $H$.The TPQ state is constructed by imaginary-time evolving a nonlocal, maximally entangled crosscap state $|\mathcal{C}\rangle$ [9, 10, 11, 12, 13, 14, 15, 16, 17] as $|\Psi_{\beta}\rangle\propto e^{-\beta H/4}|\mathcal{C}\rangle$ [19, 20, 18].
@@
The resulting dynamics reveal a distinctive, highly coherent route to equilibrium. Local observables reach to the thermal value after dephasing, yet the entanglement entropy displays a nonmonotonic evolution with a universal “double-plateau” structure. We establish this behavior from three complementary and exact perspectives. First, using 2D conformal field theory (CFT), we derive an analytical formula for the time-dependent entanglement entropy in terms of elliptic theta functions by evaluating vertex-operator correlators on the Klein bottle geometry. Second, we obtain an exact numerical benchmark by solving the governing matrix Riccati equation for the TPQ covariance matrix and evolving it in real time. Third, we show that the entire profile is quantitatively captured by an asymptotically exact quasiparticle picture [21, 22, 23, 24, 25], where the ballistic propagation of antipodally entangled quasiparticle pairs governs the coherent dephasing underlying the entropy double plateau. The perfect agreement among these approaches demonstrates that the phenomenon is exact and universal.
While it has been established that isolated integrable systems generally relax to non-thermal GGEs [26, 27], exceptions of true Gibbs-like equilibration are extremely rare and typically restricted to CFT or infinite-temperature cases [28, 29]. Our results provide an explicit lattice realization of a pure-state evolution that attains a canonical Gibbs ensemble, showing that equilibrium in integrable systems can emerge through a coherent, structured dephasing process rather than through chaos. This defines a new, exactly solvable pathway to quantum equilibration.
@@
Having obtained the exact covariance matrix $\Gamma_{\beta}$ for the initial state, the real-time evolution for the quench is given by the linear equation:
@@
The Quasiparticle Picture — The exact entanglement dynamics can be quantitatively understood through the quasiparticle picture [21, 22, 23, 24, 25]. In the standard quench paradigm from a low-entanglement state, the quench acts as a source of locally entangled quasiparticle pairs that propagate ballistically, causing entanglement to grow. The TPQ state, however, is a high-energy, extensively entangled state whose dynamics require a fundamentally different interpretation. Here, the dominant process is not the creation of entanglement, but the transport and eventual destruction of the pre-existing, non-local entanglement.
The initial state $|\Psi_{\beta}\rangle$ can be viewed as a sea of quasiparticle pairs with opposite momenta (k,k)(k,-k). Crucially, the entanglement is non-local: a quasiparticle with momentum $k$ at site $x$ is entangled not with its partner at $x$, but with the particle of momentum k-k at the antipodal site $x+L/2$. After the quench at $t=0$, these pairs propagate ballistically. The entanglement entropy of a subsystem $A$ decreases when both members of an entangled pair, which were previously shared between $A$ and its complement $\bar{A}$, become fully contained within $A$. This process leads to the following prediction for the entanglement dynamics:
@@
The function $s(k)$ is the effective entanglement density carried by each quasiparticle with momentum $k$. As proved in the appendix, the initial state is a BCS-like state in the thermodynamic limit
@@
The formula (10) provides a complete quantitative picture that accurately reproduces our exact analytical and numerical results. The initial plateau is the time before any antipodal pairs can fully enter the subsystem. The subsequent decrease is governed by the rate at which pairs enter, weighted by their entanglement density $s(k)$. The rise and second plateau are due to the same pairs exiting the subsystem as they traverse outside the subsystem. This quasiparticle picture thus provides a powerful physical interpretation for the exact thermalization observed in the system. Figure 2 shows a comparison between the exact numerical results and the quasi-particle prediction for the entanglement entropy.
Mechanism of Coherent Equilibration — We now elucidate the microscopic mechanism underlying the observed equilibration. The process is not driven by chaos but by the dephasing of the initial state’s quantum coherences, which can be described exactly in our model. After dephasing, the stationary state coincides with a GGE that, for this particular initial condition, reduces identically to the canonical Gibbs ensemble.
A defining feature of the TPQ state $|\Psi_{\beta}\rangle$ is its nonzero anomalous pairing amplitude,
$\phi_{k}(\beta)=\langle c_{k}c_{-k}\rangle=\frac{-i\sin(kN)}{2\cosh[E(k)\beta/2]}\,$, (15)
which signals quantum coherence between sectors of different fermion number. In the thermal Gibbs ensemble, such coherence vanishes identically. To understand how this coherence disappears, we analyze the BCS representation of the TPQ state in the thermodynamic limit,
$|\Psi_{\beta}(t)\rangle=\prod_{k>0}\!\left[u_{k}(\beta)+v_{k}(\beta)e^{-2iE(k)t}c_{k}^{\dagger}c_{-k}^{\dagger}\right]\!|0\rangle\,.$ (16)
The occupations $n_{k}=\langle c_{k}^{\dagger}c_{k}\rangle$ remain constant, while the pairing amplitude oscillates as $\phi_{k}(t,\beta)=\phi_{k}(0,\beta)e^{-i2E(k)t}$. Hence, observables sensitive to this coherence, such as $\langle c_{j}c_{j+1}\rangle$, display explicit temporal oscillations,
$\langle c_{j}c_{j+1}\rangle_{t}=\frac{1}{L}\sum_{k}e^{-ik(2j+1)}\phi_{k}(0)e^{-2iE(k)t}\,$, (17)
showing that the initial state is not stationary.
At long times, these oscillations dephase, and the time average satisfies $\overline{\phi_{k}(t)}=0$ for all $E(k)\neq 0$. Consequently, $\overline{\langle c_{j}c_{j+1}\rangle_{t}}=0$, exactly matching the thermal expectation value $\mathrm{Tr}(\rho_{\mathrm{th}}c_{j}c_{j+1})=0$. In contrast, number-conserving observables such as $\langle c_{j}^{\dagger}c_{j+1}\rangle$ remain stationary and coincide with their thermal values at all times. Thus, all local observables approach the canonical values once the coherence encoded in $\phi_{k}$ has dephased. The entanglement entropy provides the macroscopic witness of this microscopic process. The nonmonotonic “double-plateau” profile directly tracks the coherent loss and reorganization of pairing correlations.
After dephasing, the remaining stationary state is fully characterized by the conserved occupations $n_{k}$. The corresponding GGE density matrix reads
$\rho_{\mathrm{GGE}}=\frac{1}{Z_{\mathrm{GGE}}}\exp\!\left[-\sum_{k}\lambda_{k}\hat{n}_{k}\right]\,$, (18)
with Lagrange multipliers fixed by $\mathrm{Tr}(\rho_{\mathrm{GGE}}\hat{n}_{k})=n_{k}=(e^{\beta E(k)}+1)^{-1}$, which yields $\lambda_{k}=\beta E(k)$. Substituting this into $\rho_{\mathrm{GGE}}$ gives
$\rho_{\mathrm{st}}=\frac{e^{-\beta H}}{Z}.$ (19)
Therefore, equilibration proceeds in two coherent stages: (i) the unitary dynamics dephase the nonconserved pairing coherences, and (ii) the remaining conserved populations define a GGE that, for this special state, coincides exactly with the Gibbs ensemble. This dephasing-based mechanism provides an explicit, solvable realization of thermalization without chaos.
Conclusion and outlooks — In this work, we study the entanglement evolution following a quench from a TPQ state and uncover a distinctive, coherent pathway to equilibrium. We provide a complete theoretical account of this process and show that the stationary state coincides with the canonical Gibbs ensemble. This apparent thermal behavior arises because the special symmetries of the initial TPQ state enforce all higher conserved charges that would otherwise lead to a GGE to vanish identically. The resulting nontrivial dynamics are the macroscopic signature of an underlying microscopic mechanism. The initial state possesses nonzero anomalous pairing correlations, $\langle c_{k}c_{-k}\rangle\neq 0$, representing quantum coherence forbidden in the final thermal state. The observed “double plateau” is the real-time fingerprint of the coherent dephasing of these pairing correlations, a process driven by the ballistic propagation of stable quasiparticles. This highlights a central finding of our work that a structured, nonmonotonic entanglement evolution can emerge as the macroscopic manifestation of microscopic quantum coherence during equilibration. Conceptually, this connects our results to other observations of intermediate-time coherent dynamics in the XX chain, such as the emergence and disappearance of short-range coherence in quenches leading to an infinite-temperature state [40]. While the initial states and final ensembles differ, both reveal that the route to equilibrium in integrable systems can be remarkably rich and structured.
@@
Our results pave the way for exploring nonequilibrium dynamics from highly structured initial states and for extending the CFT quench framework to non-orientable manifolds. The CFT analysis and exact Gaussian numerics apply directly to mutual information and entanglement negativity, while the quasiparticle picture can be adapted with minor modifications. Exploring interacting integrable models in this context would be both interesting and important for future work. This perspective can be generalized to richer scenarios, including inhomogeneous quenches such as the Möbius or sine-square deformation (SSD) [41, 42, 43, 44, 45, 46]. The experimental exploration of crosscap-type initial states and the associated entanglement dynamics would be a fascinating direction for future studies, potentially realizable in quantum-simulation platforms such as ultracold atoms or superconducting qubits.
@@ Appendix C The solution of general matrix Riccati equations
Appendix D Properties of the crosscap states and TPQ states
In this section, we will thoroughly discuss various aspects of the crosscap state and TPQ state. We first observed that in the crosscap state, a pair structure in momentum space appears in the thermodynamic limit. Then we construct a parent Hamiltonian of the crosscap state, which means the ground state of this Hamiltonian is coincide with the crosscap state. In the third part, we consider the thermodynamic limit, which simplify the Hamiltonian a lot and a BCS-type wave function of the crosscap state is obtained. Finally, we derived the Fermi-Dirac distribution can be reproduced in the TPQ state.
D.1 The pair structure in the thermodynamic limit
In this section, we will prove that the crosscap state, defined on a finite chain of size $2N$, becomes a simple BCS state of independent (k,k)(k,-k) pairs in the thermodynamic limit ($N\to\infty$). The momentum distribution of the crosscap state is obtain from the correlator $C_{ij}(0)=\frac{1}{2}\delta_{ij}$ as
$\langle\mathcal{C}|c_{k}^{\dagger}c_{q}|\mathcal{C}\rangle=\frac{1}{2N}\sum_{i,j}e^{-iki}e^{iqj}\langle\mathcal{C}|c_{i}^{\dagger}c_{j}|\mathcal{C}\rangle=\frac{1}{2N}\sum_{j}e^{i(q-k)j}\left(\frac{1}{2}\right)=\frac{1}{2}\delta_{k,q}\,.$ (S92)
Thus, the normal correlation matrix is exactly diagonal for any finite system size $N$, which implies that it is also diagonal in the limit $N\to\infty$. The equation (S92) also implies that the occupation of the crosscap state is uniform $n(k,0)=1/2$ for all $k$, which is the property of an infinite temperature system.
Then we need to derive the anomalous correlator in the momentum space. The rigorous form of the real space anomalous correlator for crosscap state is
$F_{ij}=\frac{1}{2}\sum_{l=1}^{N}\left(\delta_{i,l}\delta_{j,l+N}-\delta_{j,l}\delta_{i,l+N}\right).$ (S93)
Then we compute
$\begin{split}&\langle c_{k}c_{q}\rangle=\frac{1}{2N}\sum_{i,j=1}^{2N}e^{-iki}e^{-iqj}\left[\frac{1}{2}\sum_{l=1}^{N}\left(\delta_{i,l}\delta_{j,l+N}-\delta_{j,l}\delta_{i,l+N}\right)\right]\\ &=\frac{1}{4N}\sum_{l=1}^{N}\left(e^{-ikl}e^{-iq(l+N)}-e^{-ik(l+N)}e^{-iql}\right)\\ &=\frac{1}{4N}\left(e^{-iqN}-e^{ikN}\right)\sum_{l=1}^{N}e^{-i(k+q)l}\,.\end{split}$ (S94)
The momenta are quantized on a chain of size $2N$ with anti-periodic boundary conditions
$k=\frac{\pi}{N}\left(m+\frac{1}{2}\right)\quad\text{and}\quad q=\frac{\pi}{N}\left(p+\frac{1}{2}\right),$ (S95)
where $m,p$ are integers in the range [N,N1][-N,N-1].
The core of the proof lies in the exact evaluation and analysis of the sum over half the chain. Let $\theta=k+q$. The sum in Eq. (S94) is a standard geometric series,
$S_{N}(\theta)=\sum_{l=1}^{N}e^{-i\theta l}=e^{-i\theta}\frac{1-e^{-iN\theta}}{1-e^{-i\theta}}\,.$ (S96)
This expression is exact for any $\theta\neq 2\pi n$. It is more illuminating to rewrite this in terms of sine functions using the identity $1-e^{-i\phi}=e^{-i\phi/2}(2i\sin(\phi/2))$:
$S_{N}(\theta)=e^{-i\theta}\frac{e^{-iN\theta/2}(2i\sin(N\theta/2))}{e^{-i\theta/2}(2i\sin(\theta/2))}=e^{-i(N+1)\theta/2}\frac{\sin(N\theta/2)}{\sin(\theta/2)}\,.$ (S97)
This is the well-known Dirichlet kernel. Our exact finite-$N$ correlator is therefore
$\langle c_{k}c_{q}\rangle=\frac{1}{4N}\left(e^{-iqN}-e^{ikN}\right)\left(e^{-i(N+1)(k+q)/2}\frac{\sin(N(k+q)/2)}{\sin((k+q)/2)}\right)\,.$ (S98)
We now analyze this exact expression as $N\to\infty$. Let’s consider the anti-diagonal case (q=kq=-k) first. In this case, the angle $\theta=k-k=0$. The Dirichlet kernel form is indeterminate ($0/0$), so we must return to the original sum
$S_{N}(0)=\sum_{l=1}^{N}e^{0}=N\,.$ (S99)
Now, we substitute this result back into the full expression for the correlator
$\langle c_{k}c_{-k}\rangle=\frac{1}{4N}\left(e^{ikN}-e^{-ikN}\right)N=\frac{i}{2}\sin(kN)\,.$ (S100)
Substituting the quantized momentum $k=\frac{\pi}{N}(m+1/2)$, we get
$\langle c_{k}c_{-k}\rangle=\frac{i}{2}\sin\left(\pi\left(m+\frac{1}{2}\right)\right)=\frac{i}{2}(-1)^{m}\,$, (S101)
which is of order $O(1)$ and does not vanish as $N\to\infty$. The correlation between a mode $k$ and its partner k-k persists and is finite in the thermodynamic limit.
Now consider the off-diagonal case ($q\neq-k$). In this case, the angle $\theta=k+q\neq 0$. We can use the Dirichlet kernel form. Let’s analyze the magnitude of the full expression,
$|\langle c_{k}c_{q}\rangle|=\left|\frac{1}{4N}\right|\cdot\left|e^{-iqN}-e^{ikN}\right|\cdot\left|e^{-i(N+1)\theta/2}\right|\cdot\left|\frac{\sin(N\theta/2)}{\sin(\theta/2)}\right|\leq\frac{1}{4N}\cdot 2\cdot 1\cdot\frac{1}{|\sin((k+q)/2)|}\,.$ (S102)
For any fixed $k$ and $q$ (with $q\neq-k$), the denominator $|\sin((k+q)/2)|$ is a finite number, and the entire expression is bounded by a constant times $1/N$. Therefore, in the thermodynamic limit, we have
$\lim_{N\to\infty}|\langle c_{k}c_{q}\rangle|=0\quad(\text{for }q\neq-k)\,.$ (S103)
The correlation between a mode $k$ and any other mode $q\neq-k$ vanishes in the thermodynamic limit.
In summary, we have formally proven that the momentum-space correlation matrices of the crosscap state have the following asymptotic behavior. For normal correlator, we have
$\lim_{N\to\infty}\langle c_{k}^{\dagger}c_{q}\rangle=\frac{1}{2}\delta_{k,q}\,$, (S104)
and for the anomalous correlator, we have
$\lim_{N\to\infty}\langle c_{k}c_{q}\rangle=\left(\frac{i}{2}\sin(kN)\right)\delta_{k,-q}\,.$ (S105)
The correlation matrices become exactly block-diagonal in the thermodynamic limit, which implies that the crosscap state asymptotically becomes a simple BCS state of independent (k,k)(k,-k) pairs.
D.2 The parent Hamiltonian of the crosscap states
Although the analysis in the previous section is enough to determine the final form of the crosscap state as an BCS state in the thermodynamic limit. In this section, we will try another method, which regard the crosscap state as the ground state of an special Hamiltonian– the parent Hamiltonian of the crosscap state. We find the method is inspiring and interesting in its own right.
We begin with the fundamental definition of the crosscap state in the real-space basis of fermion operators $c_{j}^{\dagger}$ on a chain of $2N$ sites:

### v3

@@ Abstract
We present an exact solution of entanglement entropy for the real-time dynamics following a quench from a thermal pure quantum (TPQ) state in an free fermion system. Contrast to the usual linear growth and saturate behavior, the entanglement entropy exhibits a characteristic double-plateau structure. We obtain this behavior exactly using three complementary approaches: two-dimensional (2D) conformal field theory (CFT) on the Klein bottle, an exact numerical evolution based on the matrix Riccati equation, and an asymptotically exact quasiparticle picture.
Understanding the non-equilibrium dynamics of isolated quantum systems is a central challenge in statistical mechanics and quantum gravity [1, 2, 3, 4, 5]. Usually, analytical and numerical studies rely on lowly entangled initial states (e.g., ground states or short-range product states), from which the celebrated linear growth and subsequent saturation of entanglement emerge [6, 7, 8]. In contrast, thermal pure quantum (TPQ) states usually construct from random state, already have volume-law entanglement and locally indistinguishable from Gibbs ensembles [9]. The entanglement dynamics from such initial states often involves exponential computational complexity, making exactly solvable models for this process very rare. Distinct from the traditional random-state-based TPQ states used for numerical calculations, recent literature has introduced highly structured TPQ state [10, 11] that have subsequently been applied in holography to elegantly capture the microstates of single-sided black holes with smooth interiors[12, 13].
This state is constructed by imaginary-time evolving a crosscap state $|\mathcal{C}\rangle$ [14, 15, 16, 17, 18, 19, 20, 21, 22] as $|\Psi_{\beta}\rangle\propto e^{-\beta H/4}|\mathcal{C}\rangle$ [11, 12, 10]. Considering that the bipartite entanglement remains constant under chaotic quenches from crosscap states [21, 12], in this Letter, we explore the quench dynamics of the bipartite entanglement entropy in the non-chaotic XX spin chain from the state $|\Psi_{\beta}\rangle$. We provide an exact analytical solution to this problem, where both the state preparation and the time evolution are governed by the same Hamiltonian $H$.
@@
We study this problem from three complementary and exact perspectives. First, using 2D conformal field theory (CFT), we derive an analytical formula for the time-dependent entanglement entropy in terms of elliptic theta functions by evaluating vertex-operator correlators on the Klein bottle geometry. Second, we obtain an exact numerical benchmark by solving the governing matrix Riccati equation for the TPQ covariance matrix and evolving it in real time. Third, we show that the entire profile is quantitatively captured by an asymptotically exact quasiparticle picture [23, 24, 25, 26], where the ballistic propagation of antipodally entangled quasiparticle pairs governs the entropy double plateau behavior.
@@
Having obtained the exact covariance matrix $\Gamma_{\beta}$ for the initial state, the real-time evolution for the quench is given by the following equation:
@@
The Quasiparticle Picture — The exact entanglement dynamics can be quantitatively understood through the quasiparticle picture [23, 24, 25, 26]. In the standard quench paradigm from a low-entanglement state, the quench acts as a source of locally entangled quasiparticle pairs that propagate ballistically, causing entanglement to grow. The TPQ state, however, is a high-energy, extensively entangled state whose dynamics require a fundamentally different interpretation. Here, the dominant process is not the creation of entanglement, but the transport and eventual destruction of the pre-existing, non-local entanglement.
The initial state $|\Psi_{\beta}\rangle$ can be viewed as a sea of quasiparticle pairs with opposite momenta (k,k)(k,-k). Crucially, the entanglement is non-local: a quasiparticle with momentum $k$ at site $x$ is entangled not with its partner at $x$, but with the particle of momentum k-k at the antipodal site $x+L/2$. After the quench at $t=0$, these pairs propagate ballistically. The entanglement entropy of a subsystem $A$ decreases when both members of an entangled pair become fully contained within $A$. This process leads to the following prediction for the entanglement dynamics:
@@
The function $s(k)$ is the effective entanglement density carried by each quasiparticle with momentum $k$. As proved in the appendix, the initial state is equivalent to a BCS-like state
@@
The formula (10) provides a complete quantitative picture that accurately reproduces our exact analytical and numerical results. The initial plateau is the time before any antipodal pairs can fully enter the subsystem. The subsequent decrease is governed by the rate at which pairs enter, weighted by their entanglement density $s(k)$. The rise and second plateau are due to the same pairs exiting the subsystem as they traverse outside the subsystem. This quasiparticle picture thus provides a powerful physical interpretation for the non-monotonic behavior of entanglement observed in the system. Figure 2 shows a comparison between the exact numerical results and the quasi-particle prediction for the entanglement entropy.
Conclusion and outlooks — In this work, we study the entanglement evolution following a quench from a volume-law entangled TPQ state. We establish this exotic entanglement dynamics rigorously through three complementary approaches: (i) a 2D CFT derivation yielding an exact analytical formula in terms of Jacobi elliptic theta functions on the Klein bottle geometry, (ii) a highly efficient exact numerical evaluation of the restricted covariance matrix, and (iii) an asymptotically exact quasiparticle picture.
@@
Our results pave the way for exploring nonequilibrium dynamics from highly entangled initial states and for extending the CFT quench framework to non-orientable manifolds. The CFT analysis and exact Gaussian numerics apply directly to mutual information and entanglement negativity, while the quasiparticle picture can be adapted with minor modifications. Exploring interacting integrable models in this context would be both interesting and important for future work. This perspective can be generalized to richer scenarios, including inhomogeneous quenches such as the Möbius or sine-square deformation (SSD) [37, 38, 39, 40, 41, 42]. The experimental exploration of crosscap-type initial states and the associated entanglement dynamics would be a fascinating direction for future studies, potentially realizable in quantum-simulation platforms such as ultracold atoms or superconducting qubits.
@@ Appendix C The solution of general matrix Riccati equations
Appendix D Exact BCS Representation of the Crosscap State
In this section, we provide an derivation of the momentum-space BCS representation of the crosscap state. By exploiting the translational symmetry inherent in the anti-periodic boundary conditions (APBC), we show that the state is strictly momentum-diagonal for any finite system size $N$.
Consider a free fermion chain of length $L=2N$ with APBC, $c_{j+2N}=-c_{j}$. The unnormalized crosscap state is defined in real space as $|\tilde{\mathcal{C}}\rangle=\exp(\hat{Q}_{N})|0\rangle$, where the generator pairs antipodal sites:
$\hat{Q}_{N}=\sum_{j=1}^{N}c_{j}^{\dagger}c_{j+N}^{\dagger}\,.$ (S92)
To diagonalize this generator in momentum space, we introduce an auxiliary operator summing over the entire lattice:
$\hat{Q}_{2N}=\sum_{j=1}^{2N}c_{j}^{\dagger}c_{j+N}^{\dagger}=\sum_{j=1}^{N}c_{j}^{\dagger}c_{j+N}^{\dagger}+\sum_{j=N+1}^{2N}c_{j}^{\dagger}c_{j+N}^{\dagger}\,.$ (S93)
By shifting the summation index l=jNl=j-N in the second term and applying the APBC ($c_{l+2N}^{\dagger}=-c_{l}^{\dagger}$), we obtain
$\sum_{l=1}^{N}c_{l+N}^{\dagger}c_{l+2N}^{\dagger}=-\sum_{l=1}^{N}c_{l+N}^{\dagger}c_{l}^{\dagger}=\sum_{l=1}^{N}c_{l}^{\dagger}c_{l+N}^{\dagger}=\hat{Q}_{N}\,$, (S94)
where the last step utilizes the fermionic anticommutation relation $c_{a}^{\dagger}c_{b}^{\dagger}=-c_{b}^{\dagger}c_{a}^{\dagger}$. This yields $\hat{Q}_{2N}=2\hat{Q}_{N}$ or equivalently
$\hat{Q}_{N}=\frac{1}{2}\hat{Q}_{2N}=\frac{1}{2}\sum_{j=1}^{2N}c_{j}^{\dagger}c_{j+N}^{\dagger}\,.$ (S95)
We now substitute the Fourier transform $c_{j}^{\dagger}=\frac{1}{\sqrt{2N}}\sum_{k}e^{-ikj}c_{k}^{\dagger}$ into the full-lattice sum. The orthogonality of the Fourier modes over the $2N$ sites rigorously enforces momentum conservation:
$\hat{Q}_{N}$ $=\frac{1}{4N}\sum_{k,q}e^{-iqN}c_{k}^{\dagger}c_{q}^{\dagger}\sum_{j=1}^{2N}e^{-i(k+q)j}$ (S96)
$=\frac{1}{4N}\sum_{k,q}e^{-iqN}c_{k}^{\dagger}c_{q}^{\dagger}\left(2N\delta_{k,-q}\right)$
$=\frac{1}{2}\sum_{k}e^{ikN}c_{k}^{\dagger}c_{-k}^{\dagger}\,.$
Under APBC, the quantized momenta are $k=\frac{\pi}{N}(m+1/2)$ with integer $m\in[-N,N-1]$. Thus, the phase factor simplifies exactly to $e^{ikN}=e^{i\pi m}e^{i\pi/2}=i(-1)^{m}$. Because the operator is strictly momentum-diagonal for any finite $N$, the normalized initial state is an exact BCS state: